An R Chart is to be developed for use in quality control. The average of the ranges is 0.30. The sample sizes were 6 each. What would the upper control limit be? O 2.3040 o.0500 O 0.6012 O 1.0982
Answer #1

1) here for sample size 6 ; factor D4 =2.004

therefore upper contrl limit =0.3*2.004=0.6012

2)

for sample size 5 ; constant D3 =0 and D4 =2.115

hence control limit =D3*Rbar and D4*Rbar =0*0.4 and 2.115*0.4 =0 and 0.846

3)

for sample size 8 ; D3 =0.136

hence lower control limit =D3*RBar =0.6*0.136=0.082